cotpie/20.cot3pie/20.cot5pie/20.cot7pie/20.cot9pie/20

Grade:12th pass

1 Answers

Vijay Mukati
8 years ago
Dear Student,

cot5pie/20 = cotpi/4 = 1
cotpie/20 = cot(pi/2 – cot9pi/20) = tan9pi/20
cot3pi/20 = cot(pi/2 – 7pi/20) = tan7pi/20
and we know that cotx * tanx = 1.
Therefore. Final answer would be = 1.

Thanks.